A crate is given an initial push on a horizontal floor and is imparted with an initial velocity of 5.0 m/s. Because of kinetic friction, the crate slows down and comes to stop with constant deceleration over a time period of 1.1 seconds. The mass of the crate is 5.0kg. what is the magnitude of the friction force that causes the crate to come to a stop? what is the coefficient of kinetic friction? (Assume that friction is the only force slowing the crate)
Answer #1

PART 1:

Given that:

Initial speed, V = 5.0 m/s

Final speed, u = 0 m/s

Time taken to stop, t = 1.1 s

Therefore, deceleration (magnitude) of the crate:

\small a=\frac{v-u}{t}=\frac{5\;m/s-0\;m/s}{1.1\;s}

\small \therefore a=4.545\;m/s^{2}

Mass of crate, m = 5kg

Therefore Magnitude of friction force is given by:

\small F=m a=(5\;kg)\times (4.545\;m/s^{2})

\small \boldsymbol{\therefore F=22.73\;N}

-------------------------------------------------------------------------

PART 2:

Friction force is given by:

\small F=\mu_{k}mg=22.73\;N

Therefore, Coefficient of kinetic friction is given by:

\small \mu_{k}=\frac{F}{mg}=\frac{22.73\;N}{5\;kg\times 9.8\;m/s^{2}}

\small \boldsymbol{\therefore \mu_{k}=0.464}

-------------------------------------------------------------------------

(Please rate the answer if you are satisfied. In case of any queries, please reach out to me via comments)